OpenStudy (anonymous):

find the complex zeros of f(x)=x^4+15x^2+14. Write f in factored form.

OpenStudy (anonymous):

x^4+15x^2+14 =0 (x^2+1)(x^2+14)=0 x^2+1 =0 or x^2 +14 =0 x^2+1-1=0-1 or x^2 +14-14=0-14 x^2=-1 or x^2=-14 x=i or -i or 14i or -14i

OpenStudy (anonymous):

f(x)=x^4+15x^2+14 =(x-i)(x+i)(x-14i)(x+14i)

OpenStudy (anonymous):

that answer is coming up as incorrect

OpenStudy (anonymous):

(x-i)(x+i)(x-14i)(x+14i) = (x^2+1)(x^2 +14) = x^4 +15x^2+14 what's the answer given?

OpenStudy (anonymous):

i don't know. i have to solve it. it just comes up as incorrect

OpenStudy (anonymous):

What are your steps?

OpenStudy (anonymous):

they don't give me any steps its just shown like i typed it in the beginning

OpenStudy (anonymous):

Oh, i see. So, why you think "that answer is coming up as incorrect" ?

OpenStudy (anonymous):

should be isqrt(14), -isqrt(14)

OpenStudy (anonymous):

because this is a homework assignment that i'm doing on a website and the program is showing the answer in incorrect.

OpenStudy (anonymous):

try \[(x-i)(x+i)(x-i \sqrt{14})(x+i \sqrt{14})\]

OpenStudy (anonymous):

I found my mistake! x^4+15x^2+14 =0 (x^2+1)(x^2+14)=0 x^2+1 =0 or x^2 +14 =0 x^2+1-1=0-1 or x^2 +14-14=0-14 x^2=-1 or x^2=-14 x=i or -i or isqrt 14 or -isqrt 14

OpenStudy (anonymous):

sorry for making mistake

OpenStudy (anonymous):

its ok thanks tho